# DiTSE: High-Fidelity Generative Speech Enhancement via Latent Diffusion Transformers

### v1

@@ Abstract
Diffusion Models, Diffusion Transformers, Speech Enhancement, Dereverberation, Generative Models.I Introduction
@@
Real-world recorded speech is often degraded by acoustic conditions such as noise, reverberation, and equalization issues. Speech enhancement methods aim to improve such signals’ perceptual quality and intelligibility [1], with useful applications to audio content creation and telecommunications. Recent advances in deep learning have established new performance standards for speech enhancement. While existing methods can alleviate some quality issues, achieving clean high-fidelity audio remains a challenge [2].
@@
The conventional enhancement methods based on deep neural networks learn a deterministic many-to-one mapping via a regressive objective, which maps input signals that have been variously degraded from a clean signal to the corresponding ground-truth clean signal. However, the problem becomes multi-modal in challenging scenarios where degradations, such as reverberation, create ambiguity in determining the target phase of the clean signal. Therefore, the conventional methods often experience mode collapse as they predict an averaged version of the many plausible signals, resulting in artifacts and distortion in their enhanced results.
@@
Generative methods directly model the distribution of clean speech, avoiding mode collapse and enabling effective utilization of speech priors. They aim to generate a plausible version of many possible high-fidelity clean signals that preserve spoken content and speaker traits of the input degraded audio. Recently, discrete generation methods for speech enhancement, often using neural codec tokens, have shown significant improvement of audio clarity and cleanness over conventional methods. They employ language modeling [3, 4] or masked language modeling [5, 6] on tokens while conditioning on the input signal. In particular, Genhancer [5] incorporates speech features to improve speaker consistency and content accuracy. However, the token spaces limited by bit-rates often miss nuances in spoken content, thus compromising the naturalness of enhanced voices. Moreover, the costly sampling processes at inference time make the methods impractical for many real-world applications.
@@
On the other hand, diffusion-based methods operate in a continuous space (i.e., sample or latent representation), stemming from the common concept of non-equilibrium thermodynamics of learning to reverse noise back to the data distribution iteratively. Particularly, SGMSE+ [7] and CDiffuSE [8] interpolate noisy speech with Gaussian noise via the forward diffusion process, aiming to reduce both the Gaussian noise and environmental degradations during the reverse process. Alternatively, UNIVERSE [9] and UNIVERSE++ [10] model speech generation conditioning on the high-level features extracted from the input audio. Moreover, StoRM [11] leverages a diffusion process to refine the outputs from other enhancement systems via regeneration, combining benefits from both conventional and generative approaches.
@@
We propose DiTSE (Diffusion Transformer for Speech Enhancement), a high-fidelity full-bandwidth enhancer using a latent diffusion transformer model. DiTSE performs diffusion in a learnt audio latent space with 32 sampling steps, effectively reducing computational burden over previous diffusion-based approaches. Comprehensive objective and subjective evaluations show that DiTSE better maintains speaker traits and reduces content hallucination compared to prior art, while achieving state-of-the-art audio quality that, for the first time, matches real studio-quality audio from the DAPS dataset [12]. We additionally introduce the technique of post-hoc quantization to further improve audio fidelity.
II Proposed Method
@@
We present the three main components of DiTSE: (1) a variational autoencoder; (2) a diffusion transformer model; (3) a conditioner network that guides the generation process with pre-trained and learned speech features. In addition, we introduce a post-hoc process to refine the audio quality using discrete generation.
II-A Preliminaries on Diffusion Models
@@
Diffusion-based models have drawn attention due to their generation capabilities in domains such as images [13] and audio [14, 15]. Diffusion models employ a forward process that adds Gaussian noise to clean input data [16], which can be expressed in marginal , where are hyperparameters from the noise scheduler [17] at the timestep . The objective is to learn to reverse this process via a neural network. However, instead of directly trying to predict the data reconstruction , it has been empirically found that other objectives are more stable to produce high-quality outputs. In this work, we use the v-prediction reparametrization [18], where the model predicts the velocity vector, , where . For the interested reader in the formulation, we refer to [19, 18, 17].
II-B Variational Autoencoder
@@
We learn a variational autoencoder on publicly available audio data (i.e., speech, music and sound events) to encode audio signals of sample rate into latent representations of 64 dimensions at frame rate. The VAE uses the network structure of a neural audio codec - DAC [20], with the quantization layers replaced by a variational bottleneck. To ensure high-quality signal reconstruction, the VAE model also adopts the GAN setup of DAC, adversarially trained with multi-period waveform discriminators and multi-scale multi-band complex spectrogram discriminators. The learned latent space, which is compressed in representation and regularized in distribution, facilitates computationally efficient generation and high-quality decoding into audio signals.
II-C Latent Diffusion Transformer (DiT) Network
@@
While U-Net architectures commonly serve as backbones for diffusion-based speech enhancement methods [21, 7], recent advancements in audio generation [22, 23] have shown that Transformer-based backbones offer significant advantages in scalability and robustness. Herein, we adopt Diffusion Transformer (DiT) [24] on the latent representations from the VAE, leveraging established practices and strategies from the vision domain for training and inference. The model conditions on the timestep embedding of the current diffusion step via adaptive layer normalization following common practices. Other conditioning information (discussed below) is provided to the network either via cross-attention or via concatenation with the diffusion latent along the feature dimension. We introduce 50% probability of prefix prompting in training, where a random number of frames at the beginning of the diffusion latent is unmasked with the ground-truth clean latent. During enhancement of long audio, we provide a previously enhanced audio segment as a prompt to select a consistent generation mode across windows of windowed generation.
II-D Robust Conditioning Features
@@
Robust SSL features. Leveraging large-scale pre-training on speech data, SSL features provide strong semantic guidance for spoken content, thus reducing content hallucinations in enhancement. This work experiments with WavLM [25] and RobustDistiller [26, 27]. The DiT network cross-attends to the learned weighted sum of feature maps of the SSL model, as the features differ in temporal resolutions from the latent.
Conditioner Network. As indicated in Figure 1, a feed-forward network using Conformer [28] architecture takes in the VAE latent and transforms it to a cleaner representation that approximates the clean latent. We jointly train the Conditioner network with the main diffusion network, using L1-loss against the ground-truth clean latent. Our preliminary study shows that stacking this pre-enhanced latent with the diffusion latent can exploit the explicit temporal alignment between the two and thus works more effectively than conditioning via cross-attention. We can view the design as if the diffusion network observes two perspectives of the clean latent at each diffusion step – one corrupted by Gaussian noise, and one corrupted by acoustic conditions translated to the latent space.
@@ II-E Post-hoc Quantization
The design above achieves good content accuracy and speaker preservation across audio samples, but displays instability in acoustic quality compared to discrete generation methods (e.g., Genhancer [5]). We hypothesize that: (1) the distance in the latent space does not correlate well with the perceptual difference, and thus, small errors could lead to perceptually significant acoustic quality degradations; (2) the diffusion trajectory is sensitive to the starting noise and sampling configurations, and is prone to accumulating errors for out-of-domain data. While prior work [10] addressed perceptual issues by fine-tuning the final diffusion steps, we propose to leverage discrete generation methods post-hoc for their benefits in audio fidelity. Specifically, we decode the predicted clean VAE latent into audio and use the discrete neural codec DAC [20] to reconstruct the audio, eliminating audible artifacts and regularizing harmonics of the speech from the latent diffusion generation.
III Experiments
III-A Experiment Setup
@@
Training Datasets. Our models are trained on publicly available datasets that simulate real-world acoustic conditions. Note that none of these datasets are used for evaluation. For clean speech, we use LibriTTS-R [29], a quality-improved 585-hour dataset, further upsampled to sample rate via bandwidth extension [30]. Noise samples are uniformly sampled from three datasets: (1) SFS-Static-Dataset [31], (2) TAU Urban Audio-Visual Scenes 2021 dataset [32], and (3) the DNS Challenge [33]. Reverberation is simulated using room impulse responses (RIRs) from the OpenSLR28 dataset [34], MIT IR Survey [35], and EchoThief [36]. To generate the pairs of synthetic training data, we employ a data simulation and augmentation procedure based on previous work [5, 37]. A speech signal is first convolved with a RIR and then mixed with one or two noise samples at a signal-to-noise ratio (SNR) uniformly sampled between -10 dB and 20 dB. The signal also undergoes random equalization and bandwidth limitation.
@@
Hyperparameters and Hardware. The DiT architecture consists of 12 Transformer layers with 8 attention heads. We evaluate two configurations: Base (335M parameters, 1024 channels, conditioned on WavLM features) and Tiny (112M parameters, 512 channels, conditioned on RobustDistiller features [26, 27], distilled from WavLM). Sinusoidal positional encodings are used across all attention layers. The conditioner branch uses a Conformer model of 12 layers, each with an input size, feed-forward dimension, and kernel size of 256, 1024, and 31, respectively. We apply Classifier-Free Guidance [38] during training with a probability of 10% to encourage the model to effectively leverage each conditioning feature. We train our models on eight NVIDIA A100 GPUs for 400k steps using a batch size of 128 5 seconds together with the AdamW optimizer [39]. The learning rate is linearly warmed up from 0 to over the first 10% training steps before decaying to in subsequent steps. We chose the cosine noise scheduler over others after qualitative comparison of the generated results. We adopt steps at inference as additional steps yield minimal improvements, and use dpmpp-3m-sde sampler from k-diffusion111https://github.com/crowsonkb/k-diffusion library.
@@
| Content [WER ] | Speaker [MOS ] | Quality [MOS ] | |||||||
| DAPS (16 kHz) | VBD (16 kHz) | EARS (16 kHz) | VBD (48 kHz) | EARS (48 kHz) | DAPS (48 kHz) | AQECC (16 kHz) | DEMO (16 kHz) | VBD (48 kHz) | |
| Input | 5.03 | 5.70 | 15.08 | 4.84 0.02 | 4.89 0.02 | 1.74 0.04 | 2.74 0.06 | 1.93 0.05 | 1.84 0.05 |
| Clean | — | — | — | 4.41 0.04 | 4.18 0.05 | 4.30 0.04 | — | — | 3.87 0.05 |
| VAE Reconst. | 0.87 | 1.23 | 3.34 | 4.28 0.04 | 4.10 0.05 | 4.25 0.04 | — | — | 3.54 0.05 |
| HiFi-GAN-2 [37] | 9.16 | 6.63 | 23.67 | 4.09 0.05 | 3.16 0.07 | 3.63 0.05 | 3.70 0.05 | 3.33 0.06 | 3.77 0.05 |
| Miipher [40] | 7.64 | 12.02 | 32.32 | 3.24 0.06 | 2.68 0.06 | 3.40 0.05 | 3.53 0.05 | 3.15 0.05 | 3.42 0.06 |
| Genhancer [5] | 6.04 | 6.47 | 26.33 | 4.03 0.05 | 3.27 0.07 | 4.08 0.05 | 3.89 0.04 | 3.78 0.05 | 4.03 0.05 |
| SGMSE+ [7] | 6.69 | 10.32 | 33.60 | 3.87 0.06 | 3.03 0.06 | 3.51 0.05 | 3.71 0.04 | 3.33 0.05 | 3.59 0.06 |
| StoRM [11] | 11.27 | 9.86 | 45.87 | 3.54 0.06 | 2.75 0.06 | 2.62 0.05 | 3.61 0.05 | 2.53 0.06 | 2.94 0.05 |
| (Ours) DiTSE-Tiny | 8.58 | 9.55 | 38.21 | 3.89 0.06 | 2.92 0.07 | 4.01 0.04 | 3.83 0.05 | 3.30 0.05 | 3.63 0.05 |
| (Ours) DiTSE-Base | 3.56 | 4.93 | 26.42 | 4.20 0.04 | 3.45 0.06 | 4.34 0.04 | 3.88 0.05 | 3.61 0.05 | 3.91 0.05 |
| (Ours) DiTSE-Base + Post | 3.71 | 5.39 | 26.72 | 4.27 0.04 | 3.50 0.06 | 4.32 0.04 | 3.97 0.04 | 3.77 0.05 | 4.00 0.05 |
III-B Ablation Study
@@
We investigate the design choices of the proposed DiTSE model on the DAPS dataset [12], which consists of pairs of studio-quality speech recordings and their degraded versions under real-world acoustic conditions. Our analysis utilizes a subset of DAPS as used by prior work [5] – one female voice (f10) and one male voice (m10) reading a 2-minute script (script5). Table I presents the objective scores of Word Error Rate (WER)222Calculated between the enhanced audio and the clean reference audio transcribed by Whisper-Large-v3-v20240930 (https://github.com/openai/whisper), WideBand Perceptual Evaluation of Speech Quality (WB-PESQ), Extended Short-Time Objective Intelligibility (ESTOI), and the non-intrusive metric DNSMOS [41]. Given the stochastic nature of generative approaches, we perform inference five times for each model using different random seeds and report the statistics across all the runs.
@@
Our Baseline approach, a DiT model without concatenating timestep embeddings or a Conditioner network, achieved notable content preservation. Incorporating additional timestep embeddings at the input improves acoustic quality. Introducing a Conditioner network further improves WER, with the masking formulation (i.e., predicting a non-negative mask) showing slightly better stability than the mapping formulation (i.e., directly predicting the features) in the latent space. Meanwhile, the smaller model, with 4.5 times fewer parameters, decreases in content accuracy but retains reasonable overall quality. Lastly, applying the post-hoc quantization to the enhanced result from our full-sized model improves its perceived audio quality, but compromises slightly content fidelity.
III-C Three-axis Evaluations

### v2

@@ Abstract
Index Terms— Diffusion Models, Diffusion Transformers, Speech Enhancement, Dereverberation, Generative Models.
1 Introduction
@@
Real-world recorded speech is often degraded by acoustic conditions such as noise, reverberation, and equalization issues. Speech enhancement (SE) methods aim to improve such signals’ perceptual quality and intelligibility, with useful applications to audio content creation and telecommunications. Recent advances in deep learning have established new performance standards for speech enhancement. While existing methods can alleviate some quality issues, achieving clean high-fidelity audio remains a challenge [1].
@@
The conventional enhancement methods based on deep neural networks learn a deterministic many-to-one mapping via a regressive objective, which maps input signals that have been variously degraded from a clean signal to the corresponding ground-truth clean signal. However, the problem becomes ill-posed in challenging scenarios where degradations, such as reverberation, create ambiguity in determining the target phase of the clean signal. Therefore, the conventional methods often experience a regression-to-the-mean behavior, resulting in artifacts and distortion in their enhanced results.
@@
Generative methods directly model the distribution of clean speech, avoiding mode collapse and enabling effective utilization of speech priors. They aim to generate a plausible version of many possible high-fidelity clean signals. Recently, discrete token generation methods [2, 3] have shown improvement in audio clarity and cleanness. Particularly, Genhancer [4] incorporates speech features to improve speaker consistency and content accuracy. However, the token space often misses nuances in spoken content, and costly inference sampling limits real-world usability.
@@
On the other hand, diffusion-based methods operate in a continuous space (i.e., sample or latent representation), stemming from the common concept of non-equilibrium thermodynamics of learning to reverse noise back to the data distribution iteratively. Particularly, SGMSE+ [5] and CDiffuSE [6] interpolate noisy speech with Gaussian noise via the forward diffusion process, aiming to reduce both the Gaussian noise and environmental degradations during the reverse process. Alternatively, UNIVERSE [7] and UNIVERSE++ [8] model speech generation conditioning on the high-level features extracted from the input audio. Moreover, StoRM [9] leverages a diffusion process to refine the outputs from other enhancement systems via regeneration, combining benefits from both conventional and generative approaches. To enhance content fidelity and speaker similarity, several approaches inject speech features as conditioning signals within the diffusion process, as demonstrated by uSEE [10]. In addition, latent diffusion models have emerged as promising alternatives, offering improved efficiency and acoustic quality [11, 12]. Among these, a concomitantly released work on speech restoration, Hi-ResLDM [11], is most closely related to our work. It adopts a two-stage architecture: first, a pretrained and frozen enhancement model recovers the degraded version of the signal; then, a latent diffusion model further refines it to improve perceptual quality.
@@
In this work, we propose DiTSE (Diffusion Transformer for Speech Enhancement), a high-fidelity, full-bandwidth model that improves speaker preservation and reduces content hallucination over prior methods, while achieving state-of-the-art audio quality that, for the first time, matches real studio quality from DAPS [13]. DiTSE performs diffusion in the latent space of a variational autoencoder, along with a robust conditioning mechanism on speech features for content and speaker guidance. By leveraging Latent Diffusion Transformers, our method demonstrates strong performance and scalability. We include comprehensive objective and subjective evaluations on audio quality, content accuracy, and speaker consistency.
2 Proposed Method
@@
DiTSE comprises three main components: (1) a variational autoencoder; (2) a diffusion transformer; and (3) a conditioner network that guides generation using both pre-trained and learned speech features. Additionally (and optional), a post-hoc refinement step improves audio quality via discrete token generation.
2.1 Preliminaries on Diffusion Models
@@
Diffusion-based models have drawn attention due to their generation capabilities [14]. Diffusion models employ a forward process that adds Gaussian noise to clean input data, which can be expressed in marginal , where are hyperparameters from the noise scheduler [15] at the timestep . The objective is to learn to reverse this process via a neural network. However, instead of directly trying to predict the data reconstruction , it has been empirically found that other objectives are more stable to produce high-quality outputs. In this work, we use the v-prediction reparametrization [16], where the model predicts the velocity vector, , where . For the interested reader in the formulation, we refer to [15].
2.2 Variational Autoencoder
@@
We train a variational autoencoder (VAE) on publicly available audio data (speech, music, and sound events) to produce 64-dimensional latent representations at a frame rate. The VAE follows the DAC [17] architecture, replacing its quantization layers with a variational bottleneck. To ensure high-fidelity reconstruction, it adopts DAC’s GAN framework, including multi-period waveform and multi-scale multi-band spectrogram discriminators. The resulting latent space is both compressed and regularized, enabling efficient generation and high-quality audio synthesis.
2.3 Latent Diffusion Transformer (DiT) Network
@@
While U-Net architectures commonly serve as backbones for diffusion-based speech enhancement methods [18, 5], recent advancements in audio generation [19] have shown that Transformer-based backbones offer significant advantages in scalability and robustness. Herein, we adopt Diffusion Transformer (DiT) [20] on the latent representations from the VAE, leveraging established practices and strategies from the vision domain for training and inference. The model conditions on the timestep embedding of the current diffusion step via adaptive layer normalization, following common practices. Other conditioning information (discussed below) is provided to the network either via cross-attention or via concatenation with the diffusion latent along the feature dimension. During training, we apply prefix prompting with a 50% probability, replacing up to the first 50% of diffusion latent frames with the corresponding ground-truth clean latent. During enhancement of long audio, we provide a previously enhanced audio segment as a prompt to ensure consistency of generation across windows of windowed generation.
2.4 Robust Conditioning Features
@@
Robust SSL features. Leveraging large-scale pre-training on speech data, SSL features provide strong semantic guidance for spoken content, thus reducing content hallucinations in enhancement. This work experiments with WavLM [21] and RobustDistiller [22]. The DiT network cross-attends to the learned weighted sum of feature maps of the SSL model, as the features differ in temporal resolutions from the VAE latent.
Conditioner Network. As indicated in Figure 1, a feed-forward network using Conformer [23] architecture takes in the VAE latent and transforms it to a cleaner representation that approximates the clean latent. We jointly train the Conditioner network with the main diffusion network, using L1-loss against the ground-truth clean latent alongside the diffusion loss. Our preliminary study shows that stacking this pre-enhanced latent with the diffusion latent can exploit the explicit temporal alignment between the two and thus works more effectively than conditioning via cross-attention. We can view the design as if the diffusion network observes two perspectives of the clean latent at each diffusion step to make its denoising decision – one corrupted by Gaussian noise, and one corrupted by acoustic conditions translated to the latent space. In addition, we experimented with masking and mapping formulations on the Conditioner branch, where it either outputs a non-negative mask to be multiplied with the input latent or directly outputs the target latent.
@@ 2.5 Post-hoc Quantization
While the design above achieves good content accuracy and speaker preservation across audio samples, it exhibits instability in acoustic quality compared to discrete token generation methods. We hypothesize two main causes: (1) a weak correlation between latent space distance and perceptual quality, making small numerical errors perceptually significant; and (2) sensitivity of the diffusion process to noise and sampling, leading to error accumulation over the trajectory for out-of-domain acoustic scenarios. While prior work [8] addressed perceptual issues by fine-tuning the final diffusion steps, we propose to leverage discrete generation methods post-hoc for their benefits in audio fidelity. Specifically, we decode the enhanced VAE latent into audio and use the discrete neural codec DAC [17] model to reconstruct the audio, eliminating audible artifacts and regularizing harmonics of the speech from the latent diffusion generation.
3 Experiments
3.1 Experiment Setup
@@
Training Datasets. Our models are trained on publicly available datasets that simulate real-world acoustic conditions. Note that none of these datasets are used for evaluation. For clean speech, we use LibriTTS-R [24], a quality-improved 585-hour dataset, further upsampled to sample rate via bandwidth extension [25]. Noise samples are uniformly sampled from three datasets: (1) SFS-Static-Dataset [26], (2) TAU Urban Audio-Visual Scenes 2021 dataset [27], and (3) the DNS Challenge [28]. Reverberation is simulated using room impulse responses (RIRs) from the OpenSLR28 dataset [29], MIT IR Survey [30], and EchoThief [31]. To generate the pairs of synthetic training data, we employ a data simulation and augmentation procedure based on previous work [4]. A speech signal is first convolved with a RIR and then mixed with one or two noise samples at a signal-to-noise ratio (SNR) uniformly sampled between -10 dB and 20 dB. The signal also undergoes random equalization and bandwidth limitation.
@@
Hyperparameters and Hardware. The DiT architecture consists of 12 Transformer layers with 8 attention heads (335M parameters, 1024 channels, conditioned on WavLM features). We also do ablation with a small version (112M parameters, 512 channels, conditioned on RobustDistiller features [22], distilled from WavLM). Sinusoidal positional encodings are used across all attention layers. The conditioner branch uses a Conformer model of 12 layers, each with an input size, feed-forward dimension, and kernel size of 256, 1024, and 31, respectively. We apply Classifier-Free Guidance [32] during training with a probability of 10% to encourage the model to effectively leverage each conditioning feature. We train our models on eight NVIDIA A100 GPUs for 400k steps using a batch size of 128 5 seconds together with the AdamW optimizer. The learning rate is linearly warmed up from 0 to over the first 10% training steps before decaying to in subsequent steps. We chose the cosine noise scheduler after qualitatively comparing the generated results to ones from other noise schedulers. We adopt steps at inference as additional steps yield minimal improvements, and use dpmpp-3m-sde sampler from k-diffusion111https://github.com/crowsonkb/k-diffusion library.
@@
| (-) Scaling Down | 8.49 0.45 | 2.12 0.01 | 81.92 0.11 | 3.31 0.01 |
| Content [WER ] | Speaker [MOS ] | Quality [MOS ] | ||||||||
| SR (kHz) | DAPS (16 kHz) | VBD (16 kHz) | EARS (16 kHz) | VBD (48 kHz) | EARS (48 kHz) | DAPS (48 kHz) | AQECC (16 kHz) | DEMO (16 kHz) | VBD (48 kHz) | |
| Input | 48 | 5.03 | 5.70 | 15.08 | 4.84 0.02 | 4.89 0.02 | 1.74 0.04 | 2.74 0.06 | 1.93 0.05 | 1.84 0.05 |
| Clean | 48 | — | — | — | 4.41 0.04 | 4.18 0.05 | 4.30 0.04 | — | — | 3.87 0.05 |
| VAE Reconst. | 48 | 0.87 | 1.23 | 3.34 | 4.28 0.04 | 4.10 0.05 | 4.25 0.04 | — | — | 3.54 0.05 |
| HiFi-GAN-2 [33] | 48 | 9.16 | 6.63 | 23.67 | 4.09 0.05 | 3.16 0.07 | 3.63 0.05 | 3.70 0.05 | 3.33 0.06 | 3.77 0.05 |
| Miipher [34] | 22.1 | 7.64 | 12.02 | 32.32 | 3.24 0.06 | 2.68 0.06 | 3.40 0.05 | 3.53 0.05 | 3.15 0.05 | 3.42 0.06 |
| Genhancer [4] | 44.1 | 6.04 | 6.47 | 26.33 | 4.03 0.05 | 3.27 0.07 | 4.08 0.05 | 3.89 0.04 | 3.78 0.05 | 4.03 0.05 |
| SGMSE+ [5] | 48 | 6.69 | 10.32 | 33.60 | 3.87 0.06 | 3.03 0.06 | 3.51 0.05 | 3.71 0.04 | 3.33 0.05 | 3.59 0.06 |
| StoRM [9] | 16 | 11.27 | 9.86 | 45.87 | 3.54 0.06 | 2.75 0.06 | 2.62 0.05 | 3.61 0.05 | 2.53 0.06 | 2.94 0.05 |
| (Ours) DiTSE | 48 | 3.56 | 4.93 | 26.42 | 4.20 0.04 | 3.45 0.06 | 4.34 0.04 | 3.88 0.05 | 3.61 0.05 | 3.91 0.05 |
| (Ours) DiTSE + Post | 48 | 3.71 | 5.39 | 26.72 | 4.27 0.04 | 3.50 0.06 | 4.32 0.04 | 3.97 0.04 | 3.77 0.05 | 4.00 0.05 |
3.2 Ablation Study
@@
We investigate the design choices of the proposed DiTSE model on the DAPS dataset [13], which consists of pairs of studio-quality speech recordings and their degraded versions under real-world acoustic conditions. Our analysis utilizes a subset of DAPS as used by prior work [4] – one female voice (f10) and one male voice (m10) reading a 2-minute script (script5). Table 1 presents the objective scores of Word Error Rate (WER)222Calculated between the enhanced audio and the clean reference audio transcribed by Whisper-Large-v3-v20240930 (https://github.com/openai/whisper), WideBand Perceptual Evaluation of Speech Quality (WB-PESQ), Extended Short-Time Objective Intelligibility (ESTOI), and the non-intrusive metric DNSMOS [35]. Given the stochastic nature of generative approaches, we perform inference five times for each model using different random seeds and report the statistics across all the runs.
@@
Our Baseline approach, a DiT model without auxiliary timestep embeddings or Conditioner network, achieved notable content preservation. Incorporating additional timestep embeddings at the input improves acoustic quality. Introducing Conditioner network further improves WER, and the masking formulation shows better stability than the mapping formulation in the latent space. Lastly, applying the post-hoc quantization to the enhanced audio from our full-sized model improves its perceived quality, but compromises content fidelity slightly.
While our model is larger than the baselines, the comparison remains valid and highlights the strengths of our approach. We also evaluated a 4.5 times smaller version, which showed reduced content accuracy but maintained overall quality comparable to diffusion-based baselines. Although still about 2 times larger than SGMSE+ in terms of parameters, our model benefits from a DiT-based architecture that diffuses in the VAE’s latent space, allowing more efficient training (e.g., four times larger batch sizes) and significantly better acoustic quality with scaling, an advantage that is less evident in U-Net-based generative models on time-frequency representations.
3.3 Three-axis Evaluations